OpenStudy (anonymous):

each interior angle of a regular polygon has a measure of 180 degrees. how many sides does it have?

OpenStudy (gorv):

u know when each angle become 180 degree

OpenStudy (gorv):

is it posibble to have each angle 180 degree

OpenStudy (gorv):

each angle=180-360/n here n = no of side

OpenStudy (anonymous):

Sum of interior angles = 180(n - 2)

OpenStudy (gorv):

it will be a circle